OpenStudy (wwe123):

find value of x......... 4^x-3^[x-(1/2)] = 3^[x+(1/2)]-2^(2x-1)

OpenStudy (anonymous):

3^(x-1/2)=a 4^x=b b-a=3a+1/2b 4a=b/2 b=8a......

OpenStudy (wwe123):

demonstrate more

OpenStudy (anonymous):

3^(x-1/2)=a 4^x=b b-a=3a+1/2b 4a=b/2 b=8a......

OpenStudy (wwe123):

how 3^[x+(1/2)]= 3a ??

OpenStudy (wwe123):

& 2^[2x-1]=1/2 b ??

OpenStudy (callisto):

As you see in @mahmit2012 's post, Let \(3^{(x-\frac{1}{2})}=a\) \[3^{(x+\frac{1}{2})}= 3^{(x-\frac{1}{2})+1}=3^{(x-\frac{1}{2}) \times 3} = a \times 3 = 3a\]

OpenStudy (callisto):

And let \(4^x=b\), That is \(2^{2x} = b\) So, \[2^{2x-1}=2^{2x}\times \frac{1}{2} = b \times \frac{1}{2} = \frac{b}{2}\] You can get these using laws of indices.

OpenStudy (campbell_st):

well \[4^x = 2^{2x}\] slight error in the notation \[3 \times 3^{x - \frac{1}{2}} = 3^{x - \frac{1}{2} + 1} = 3^{ x + \frac{1}{2}}\]

OpenStudy (wwe123):

thanks to callisto,mahmit2012 & all

OpenStudy (callisto):

Sorry.. Something's wrong with my first post here. It should be \[3^{(x+\frac{1}{2})}= 3^{(x-\frac{1}{2})+1}=3^{(x-\frac{1}{2})} \times 3 = a \times 3 = 3a\]

OpenStudy (anonymous):

3(x+12)=3(x−12)+1=3(x−12)×3=a×3=3a

OpenStudy (wwe123):

but how to solve for x ??

OpenStudy (callisto):

you get b=8a right now, right?

OpenStudy (wwe123):

??

OpenStudy (callisto):

Do you know how to get b=8a?

OpenStudy (callisto):

If you do, I'll go to the next step; if not, I'll have to explain it!

OpenStudy (wwe123):

yes, i know

OpenStudy (callisto):

Okay, so, now we get b=8a. Put \( a = 3^{(x-\frac{1}{2})}\) and \(4^x=b\) into the equation. What do you get?

OpenStudy (wwe123):

b=8a 4

OpenStudy (wwe123):

b=8a 4^x= 8*3[x-(1/2)]

OpenStudy (callisto):

Yes. Now, take log on both sides. Have you learnt logarithm yet?

OpenStudy (wwe123):

yes,i am to equal the base like [2^x=2^3,so x=3]

OpenStudy (wwe123):

yes,but i am trying to equal the base like [2^x=2^3,so x=3]

OpenStudy (callisto):

Hmm..Hope you know how it works.. When you solve 2^x=2^3, since they have the same base, you can equate the power (as it must be equal when they share the same base) But, say, you have 5^x = 2^3, you can use log to solve x as the following: 5^x = 2^3 Take log on both sides ln (5^x) =ln (2^3) By log property: ln (a^b) = b ln (a), so you'll get: x ln5 = ln 2^3 (note that 2^3 = 8) Divide both sides by ln5 to isolate x: x = (ln8) / (ln5) Got it?

OpenStudy (wwe123):

yes

OpenStudy (wwe123):

so how to solve 4^x= 8*3[x-(1/2)] by log

OpenStudy (callisto):

Take log for both sides like I did.

OpenStudy (wwe123):

log 2^2x =log 2^3*3^[x-(1/2)]

OpenStudy (callisto):

Not really.. Instead: \[\ln (4^x) = \ln (8\times 3^{x-\frac{1}{2}})\]Got it so far?

OpenStudy (wwe123):

and other step

OpenStudy (callisto):

Do you understand that?

OpenStudy (wwe123):

yes i understand but not able to solve further

OpenStudy (callisto):

Okay, after you get that step, you have to apply the log. properties: (1) \(\ln a^b = b\ln a\) (2) \(\ln ab = \ln a + \ln b \) Any problem so far?

OpenStudy (wwe123):

i know all the rule of log

OpenStudy (callisto):

Alright, so it's all about application here. \[\ln (4^x) = \ln (8\times 3^{x-\frac{1}{2}})\]\[\ln (4^x) = \ln (8) + \ln (3^{x-\frac{1}{2}})\]Any problems?

OpenStudy (wwe123):

yes, no problem

OpenStudy (callisto):

Apply the log. property (2). \[x\ln (4) = \ln (8) + ({x-\frac{1}{2}})\ln (3)\] Any questions?

OpenStudy (wwe123):

yes, no problem

OpenStudy (callisto):

\[x\ln (4) = \ln (8) + {x \ln3 -(\frac{1}{2}})\ln (3)\] Got it?

OpenStudy (wwe123):

yes

OpenStudy (callisto):

Now, solve the equation by simplifying the constant term and doing the addition/subtraction for like terms. Can you do it?

OpenStudy (wwe123):

sorry, more..

OpenStudy (callisto):

One more hint: Subtract xln3 from both sides: \[x\ln (4) - x \ln3= \ln (8) -(\frac{1}{2})\ln (3)+ x \ln3 - x \ln3\]\[x(\ln (4) - \ln3)= \ln (8) -(\frac{1}{2})\ln (3)\]Can you do it now?

OpenStudy (callisto):

@wwe123 Where are you stuck at?

OpenStudy (wwe123):

in find the value of x x(ln(4)−ln3)=ln(8)−(12)ln(3)

OpenStudy (wwe123):

it not 12 ,1/2

OpenStudy (callisto):

The rest is just sum manipulations. Can you tell me exactly where you don't know how to do?

OpenStudy (wwe123):

(4^x)/3 =8/(3½)

OpenStudy (callisto):

How do you get that?

OpenStudy (wwe123):

for got that plz solve x(ln(4)−ln3)=ln(8)−(1/2)ln(3)

OpenStudy (callisto):

How would you solve this equation?

OpenStudy (wwe123):

no

OpenStudy (callisto):

Have you tried to solve it?

OpenStudy (wwe123):

yes, a have a qurry this question is of quadratic equation

OpenStudy (callisto):

quadratic? Can you show me how you solve it?

OpenStudy (wwe123):

no i am not able solve it . neither by quadratic equation nor by log

OpenStudy (callisto):

just show me your attempt..

OpenStudy (wwe123):

i try your method but not able to solve it by quadratic equation

OpenStudy (wwe123):

it is quadratic equation, how we take 2 variable & solve it by log

OpenStudy (callisto):

Why is it a quadratic equation?

OpenStudy (callisto):

There is only one variable in this question, that is x.

OpenStudy (wwe123):

solve this ques ans is x=(3/2)

OpenStudy (callisto):

how do you get this? Can you show your workings?

OpenStudy (wwe123):

i know ans only becoz its give in my math book

OpenStudy (wwe123):

((((quadratic equation)))))

OpenStudy (callisto):

Phew. We did something wrong!

OpenStudy (callisto):

Let's start all over again! The question is \(4^x-3^{(x-\frac{1}{2})} = 3^{(x+\frac{1}{2})}-2^{(2x-1)}\) Let \(a =3^{(x-\frac{1}{2})}\) and \(b=4^x=2^{2x}\) \[3^{(x+\frac{1}{2})} = 3^{(x-\frac{1}{2})+1}= 3^{(x-\frac{1}{2})} \times 3 =3a\]\[2^{(2x-1)}= \frac{2^{2x}}{2}=\frac{b}{2}\] \[4^x-3^{(x-\frac{1}{2})} = 3^{(x+\frac{1}{2})}-2^{(2x-1)}\]\[b - a = 3a - \frac{b}{2}\]\[\frac{3b}{2} = 4a\]\[b =\frac{ 8a}{3}\]

OpenStudy (callisto):

Now, put \(a =3^{(x-\frac{1}{2})}\) and \(b=4^x\) into the equation. \[4^x = \frac{8}{3} (3^{(x-\frac{1}{2})})\]Take log. for both sides.\[\ln 4^x = \ln [\frac{8}{3} (3^{(x-\frac{1}{2})})]\]Use log property log(ab) = log (a) + log (b) for the right side.\[\ln 4^x = \ln (\frac{8}{3}) + \ln (3^{(x-\frac{1}{2})})\]Apply the log property log (a^b) = b log(a) for both sides.\[x \ln 4 = \ln (\frac{8}{3}) + (x-\frac{1}{2})\ln (3)\]\[x \ln 4 = \ln (\frac{8}{3}) + x\ln (3)-\frac{1}{2}\ln (3)\]

OpenStudy (wwe123):

wait

OpenStudy (wwe123):

euraka ,i done it

OpenStudy (wwe123):

do not use log

OpenStudy (wwe123):

its quadratic equation

OpenStudy (wwe123):

see this

OpenStudy (wwe123):

b=8a/3

OpenStudy (wwe123):

4^x = (8/3)*3^[x-(1/2)] 2^2x = 8^[x-(1/2)] 2^2x = 2^3[x-(1/2)] 2x=3[x-(1/2)] 2x=3x-(3/2) 2x=(6x-3)/2 4x=6x-3 2x-3=0 x=3/2 and answer is (3/2)

OpenStudy (callisto):

Sorry, would you mind explaining the step 2^2x = 8^[x-(1/2)], how do you get this?

OpenStudy (callisto):

\[2^{2x} = \frac{8}{3} \times 3^{(x-\frac{1}{2})}\]\[2^{2x} = 8\times 3^{(x-\frac{1}{2})-1}\]\[2^{2x} = 2^3\times 3^{(x-\frac{3}{2})}\]Divide both sides by 2^3 \[\frac{2^{2x}}{2^3} = \frac{2^3\times 3^{(x-\frac{3}{2})}}{2^3}\]\[2^{2x-3} = 3^{(x-\frac{3}{2})}\] You still need to take log to solve it!